# Learning Network Representations with Disentangled Graph Auto-Encoder

### v1

@@ 6.1 Link prediction
Table 2: Results for Link Prediction of citation networks.(*** denotes dataset without features, i.e., $X=\rm{I}$). Higher is better.
Model Cora CiteSeer PubMed Cora* CiteSeer* PubMed*
Auc AP Auc AP Auc AP Auc AP Auc AP Auc AP
SC - - - - - - $89.4\pm 1.42$ $89.8\pm 1.44$ $90.1\pm 1.2$ $89.2\pm 1.31$ $90.1\pm 0.38$ $87.5\pm 0.65$
DW - - - - - - $85\pm 1.29$ $86.4\pm 0.91$ $89\pm 0.99$ $90.6\pm 0.72$ $91.7\pm 0.11$ $92.2\pm 0.22$
node2vec - - - - - - $85.7\pm 1.09$ $88.7\pm 1.27$ $90.3\pm 1.22$ $92.9\pm 0.74$ $93.9\pm 0.13$ $94.6\pm 0.15$
GAE $91\pm 1.02$ $91.9\pm 0.89$ $90.1\pm 1.28$ $92\pm 0.96$ $96.6\pm 0.06$ $96.9\pm 0.04$ $83.8\pm 1.37$ $88.4\pm 0.74$ $76\pm 1.11$ $82.8\pm 0.56$ $81.2\pm 0.33$ $87.3\pm 0.27$
LGAE $93.7\pm 0.38$ $94.5\pm 0.26$ $94.9\pm 0.88$ $95.4\pm 0.9$ $97.6\pm 0.08$ $97.7\pm 0.12$ $85\pm 1.21$ $89.4\pm 0.92$ $79.5\pm 1$ $84.9\pm 0.72$ $83.5\pm 0.45$ $88.5\pm 0.4$
ARGA $93.1\pm 0.34$ $93.2\pm 0.44$ $93.9\pm 0.63$ $94.3\pm 0.65$ $91.1\pm 1.04$ $90.1\pm 1.08$ $76.4\pm 2.62$ $78\pm 2.11$ $71.5\pm 0.87$ $76.4\pm 0.76$ $79\pm 0.4$ $84.9\pm 0.38$
GNAE $95\pm 0.34$ $95.6\pm 0.25$ $95.1\pm 0.48$ $96.1\pm 0.4$ $97.5\pm 0.11$ $97.4\pm 0.09$ $87\pm 1.26$ $91.1\pm 0.81$ $79.2\pm 1.2$ $85.2\pm 0.94$ $84.4\pm 0.13$ $89.8\pm 0.13$
DGA $\boldsymbol{99.5\pm 0.11}$ $\boldsymbol{99.6\pm 0.03}$ $\boldsymbol{99.7\pm 0.17}$ $\boldsymbol{99.7\pm 0.12}$ $\boldsymbol{99.2\pm 0.04}$ $\boldsymbol{99.2\pm 0.05}$ $\boldsymbol{99.4\pm 0.13}$ $\boldsymbol{99.5\pm 0.11}$ $\boldsymbol{99.1\pm 0.07}$ $\boldsymbol{99.2\pm 0.05}$ $\boldsymbol{98.7\pm 0.13}$ $\boldsymbol{98.8\pm 0.11}$
VGAE $90.9\pm 0.4$ $92.5\pm 0.28$ $89.2\pm 0.99$ $91.3\pm 0.55$ $95.7\pm 0.21$ $96\pm 0.2$ $84.8\pm 1.26$ $88.5\pm 0.85$ $76.4\pm 0.92$ $82.3\pm 0.47$ $82.3\pm 0.43$ $87.8\pm 0.29$
LVGAE $94.7\pm 0.67$ $94.9\pm 0.82$ $94.9\pm 0.36$ $95.4\pm 0.35$ $97.6\pm 0.15$ $97.6\pm 0.23$ $84.8\pm 1.3$ $88.8\pm 0.98$ $79.1\pm 1.2$ $84.2\pm 0.91$ $84.1\pm 0.6$ $88.9\pm 0.42$
ARVGA $91.1\pm 0.7$ $91.2\pm 0.74$ $91.9\pm 1.21$ $92.5\pm 1.13$ $95.4\pm 0.2$ $95.9\pm 0.1$ $68.9\pm 4.07$ $72.6\pm 3.59$ $71.4\pm 1.74$ $75.6\pm 2.23$ $84\pm 0.57$ $87.9\pm 0.48$
SIGVAE $91.5\pm 0.68$ $92.6\pm 0.71$ $89.5\pm 0.82$ $91.4\pm 0.57$ $85.1\pm 0.8$ $88.5\pm 0.6$ $77.6\pm 1.58$ $83.4\pm 1.46$
VGNAE $95.1\pm 1.27$ $95.6\pm 1.17$ $95.1\pm 0.54$ $96\pm 0.55$ $97.4\pm 0.15$ $97.4\pm 0.14$ $86\pm 0.97$ $90.1\pm 0.55$ $78.7\pm 1.46$ $85\pm 1.23$ $84.6\pm 0.44$ $89.7\pm 0.36$
DVGA $\boldsymbol{97.1\pm 0.21}$ $\boldsymbol{97.2\pm 0.18}$ $\boldsymbol{95.6\pm 0.05}$ $\boldsymbol{96.2\pm 0.12}$ $\boldsymbol{97.6\pm 0.11}$ $\boldsymbol{97.8\pm 0.06}$ $\boldsymbol{97\pm 0.15}$ $\boldsymbol{97.2\pm 0.21}$ $\boldsymbol{94.6\pm 0.48}$ $\boldsymbol{95.3\pm 0.37}$ $\boldsymbol{97.8\pm 0.09}$ $\boldsymbol{98\pm 0.07}$
@@
Table 3: Results for Link Prediction of synthetic graphs with varying numbers of latent factors.(*** denotes dataset without features, i.e., $X=\rm{I}$). Higher is better.
@@
(6) Experimental Results: Table 2 presents the results of link prediction on citation networks. As observed, DGA and DVGA consistently outperform other GAE-based or VGAE-based methods, which demonstrates the effectiveness of our models structure in generating more accurate predictions. This effectiveness results in substantial improvements in the performance of link prediction on real-world datasets. For instance, DVGA achieves a prediction accuracy increase of 2.1% and 0.5% when compared to the best-performing VGAE-based baselines on Cora and CiteSeer, respectively. Similarly, DGA demonstrates prediction accuracy improvements of 4.7%, 4.8%, and 1.6% in terms of AUC against the strongest GAE-Based baselines on Cora, CiteSeer, and PubMed, respectively.
The formation processes of real-world graphs are typically unobservable, posing challenges in acquiring semantic information about latent factors. To further investigate the behavior of our models, as mentioned earlier before, we generate synthetic datasets. We present link prediction results in Table 3 with varying the number of channels ($K$) for latent factors. Our results indicate that our models consistently outperform baselines when varying the number of latent factors. Particularly, as the number of factors increases from 4 to 10, the performance of GAE and VGAE steadily decreases. In contrast, our model’s performance only exhibits a slight decrease, demonstrating its insensitivity to $K$. Especially, DVGA significantly outperforms VGAE by 27.3% in terms of AUC and 36.9% in terms of AP on the synthetic graph with ten latent factors.
These results arise from the limitations of existing methods in identifying the underlying latent factors crucial for exploring graph properties, thus hindering the learning of disentangled representations. In contrast, our approach explicitly takes into account the entanglement of heterogeneous factors and designs a decoder specifically suited for this context. Consequently, in the task of link prediction, our method exhibits superior accuracy compared to other approaches across all three datasets.
6.2 Node clustering
In this section, we address the task of unsupervised node clustering in the graph. We utilize the K-means clustering algorithm on the learned node embeddings, setting $K$ according to the number of classes specified in Table 1.
Table 4: Node clustering results on Cora and CiteSeer.
@@ 6.2 Node clustering
(1) Evaluation metrics: The node labels provided in Table 1 serve as the true clustering labels for our evaluation. We adopt the approach introduced in Xia et al. (2014), where we initially use the Munkres assignment algorithm (Munkres, 1957) to match the predicted labels with the true labels. Subsequently, we employ five metrics for validating the clustering results, specifically reporting (i) accuracy (acc); (ii) precision; (iii) F1-score (F1); (iv) normalized mutual information (NMI); (v) average rand index (ARI). Consistent with previous studies such as Pan et al. (2018) and Shi et al. (2020), our focus is only on Cora and CiteSeer.
(2) Experimental Results: Table 4 shows the results of node clustering for Cora and CiteSeer. The results indicate that both DGA and DVGA exhibit a substantial enhancement across all five metrics in comparison to the remaining baseline methods. For instance, on CiteSeer, DVGA has demonstrated a 49.4% accuracy improvement compared to GNAE and a 45.8% improvement compared to DW. Additionally, it has achieved a 2.5% increase in NMI compared to LVGAE and a 78.8% increase compared to DW. On Cora, DGA has exhibited a 71.6% precision enhancement compared to VGAE and a 28.0% improvement compared to DW. Furthermore, it has shown a 0.6% increase in NMI compared to GNAE and a 27.4% improvement compared to DW. This results further proved the effectiveness and superiority of our model.
6.3 Semi-supervised node classification
The node classification task aims to predict the labels of the remaining nodes by utilizing the labels of a subset of nodes in the underlying graph. We adapt DGA and DVGA for node classification on three citation networks by adjusting the loss function to incorporate both the original model loss and a semi-supervised classification term. A hyper-parameter is introduced to control the relative importance of components in the combined objective function. We employ an additional fully-connected neural network to map the latent representations to softmax probabilities for the required number of classes. During the optimization process, all parameters are trained jointly.
Table 5: Classification accuracies (in percent). Baseline results from Kipf et al. (2016b).
@@ 6.3 Semi-supervised node classification
Experimental Results: The results of node classification are summarized in Table 5. We observe that our models outperform other competing models on Cora and PubMed. It also shows comparable performance to CiteSeer, where it closely aligns with the GCN, considered the best baseline. The competitive performance in comparison to state-of-the-art methods highlights the robust capabilities of our model, even though it is not specifically trained for this task.
6.4 Qualitative Evaluation
For a more comprehensive understanding of our proposed models, we conduct two qualitative experiments for a thorough examination in parallel with VGAE. These evaluations focus on performance of disentanglement and the informativeness learned in the embeddings.
(1) Correlation of Latent Features:
@@ 6.4 Qualitative Evaluation
We plot the correlation of the dimensions of latent features on the synthetic graph with four latent factors in Figure 3. The figure illustrates the absolute correlation values between elements within the 64-dimensional graph representations obtained from VGAE and DVGA. DVGA’s correlation plot exhibits a clear blockwise correlation pattern, specifically with four distinct diagonal blocks. This suggests that the four channels in DVGA are likely to capture mutually exclusive information. Therefore, our proposed method successfully learns disentangled node embeddings and captures the latent factors leading to connections to some degree. This will provide interpretability for the link prediction results.
(2) Visualization of Node Embeddings:
@@
Figure 4 provides a visual comparison of the learned node embeddings from DW, VGAE, ARVGA, and DVGA on Cora. Utilizing t-SNE (Van et al., 2008), we project the node representations into a two-dimensional space. It is evident that DVGA generally learns superior node embeddings, showcasing high intraclass similarity and interclass differences. These findings emphasize the effectiveness of DVGA in learning disentangled representations of graph-structured data.
6.5 Analysis
Table 6: Ablation Studies of DGA.
@@ 6.5 Analysis
Table 7: Ablation Studies of DVGA.
@@
(1) Ablation Studies: We conduct ablation studies on the essential components of our models to validate their contributions. We compare DVGA with the following three variants: (1) $w/o$ flow: i.e., setting $M=0$. (2) $w/o$ indep: i.e., setting $\lambda=0$. (3) $w/o$ disen: i.e., setting $K=1$, and in this case, our method degenerates into the traditional entangled variational graph auto-encoder. For DGA, as it lacks the flow component in its structure, we consider its similar variants $w/o$ indep and $w/o$ disen. For simplicity, we only report the results of the unsupervised link prediction task, with similar patterns observed in other tasks.
@@
The results for DGA, DVGA, and their variants are presented in Tables 6 and 7. It is evident that removing any module diminishes the predictive capability of the models across all datasets. Especially noteworthy is the substantial impact of eliminating the disentanglement mechanism, as observed in the significant reduction in model performance. For instance, the results of DVGA on Cora illustrate this effect. Therefore, we conclude that each module we designed contributes significantly to the models’ capability, and their presence collectively enhances the predictive power of the models.
(2) Hyper-Parameter Sensitivity: We explore the sensitivity of three crucial hyper-parameters. We present the results for evaluating the channel $K$ on synthetic datasets (Figure 5) and the results for all three hyperparameters on Cora (Figure 6), CiteSeer (Figure 7), and PubMed (Figure 8), with the number of channels $K$ recognized as the most significant one. Here, we present the results of DVGA under the AUC metric for link prediction. Similar findings are observed for the AP metric, as illustrated in Figures 10, 11, and 12 in Appendix.
  • A)
@@
    Number of channels $K$: We initially examine $K$ on the synthetic dataset with eight latent factors. It can be seen that the performance initially increases with an increment in $K$, reaching its peak at $K=8$, and subsequently decreases. This suggests that an optimal $K$ aligning with the ground-truth latent factors outputs the best results. It also illustrates that our model learns representations that promote disentanglement and are effective for node-related tasks. Then we investigate the impact of the channel $K$ in three citation networks. As observed previously, the performance initially increases, then reaches a peak before experiencing a notable declining. This pattern also shows that an appropriate number of channels $K$, matching the real number of latent factors can result in improved performance;
  • B)
@@
    Regularization coefficient $\lambda$: The regularization coefficient $\lambda$ can affect the model’s performance. A large $\lambda$ tends to excessively emphasize the independence between the latent factors, whereas an excessively small $\lambda$ constrains the influence of the independence regularization. Through empirical observation, we find that setting $\lambda$ to 0.01, 0.1, and 0.001 respectively leads to satisfactory results for Cora, CiteSeer, and PubMed.
  • C)
@@
    Number of disentangle layers $L$: The significance of the number of disentangle layers $L$ lies in the fact that a model with a small $L$ exhibits restricted capacity, potentially hindering its ability to integrate sufficient information from neighbors. From the graphs, we observe that for Cora, the optimal $L$ is 3; for CiteSeer, it is 5; and for PubMed, it is 7. Additionally, when $L$ is too large, there is a decline in predictive capability, suggesting that excessive message propagation can result in overfitting, thereby compromising model performance;
6.6 Convergence Behavior and Complexity Analysis

### v2

@@ VI-A Link prediction
TABLE II: Results for Link Prediction of citation networks.(* denotes dataset without features, i.e., $X=\rm{I}$). Higher is better.
Model Cora* CiteSeer* PubMed*
Auc AP Auc AP Auc AP
SC $89.4\pm 1.42$ $89.8\pm 1.44$ $90.1\pm 1.2$ $89.2\pm 1.31$ $90.1\pm 0.38$ $87.5\pm 0.65$
DW $85\pm 1.29$ $86.4\pm 0.91$ $89\pm 0.99$ $90.6\pm 0.72$ $91.7\pm 0.11$ $92.2\pm 0.22$
node2vec $85.7\pm 1.09$ $88.7\pm 1.27$ $90.3\pm 1.22$ $92.9\pm 0.74$ $93.9\pm 0.13$ $94.6\pm 0.15$
GAE $83.8\pm 1.37$ $88.4\pm 0.74$ $76\pm 1.11$ $82.8\pm 0.56$ $81.2\pm 0.33$ $87.3\pm 0.27$
LGAE $85\pm 1.21$ $89.4\pm 0.92$ $79.5\pm 1$ $84.9\pm 0.72$ $83.5\pm 0.45$ $88.5\pm 0.4$
ARGA $76.4\pm 2.62$ $78\pm 2.11$ $71.5\pm 0.87$ $76.4\pm 0.76$ $79\pm 0.4$ $84.9\pm 0.38$
GNAE $87\pm 1.26$ $91.1\pm 0.81$ $79.2\pm 1.2$ $85.2\pm 0.94$ $84.4\pm 0.13$ $89.8\pm 0.13$
DGA $\boldsymbol{99.4\pm 0.13}$ $\boldsymbol{99.5\pm 0.11}$ $\boldsymbol{99.1\pm 0.07}$ $\boldsymbol{99.2\pm 0.05}$ $\boldsymbol{98.7\pm 0.13}$ $\boldsymbol{98.8\pm 0.11}$
VGAE $84.8\pm 1.26$ $88.5\pm 0.85$ $76.4\pm 0.92$ $82.3\pm 0.47$ $82.3\pm 0.43$ $87.8\pm 0.29$
LVGAE $84.8\pm 1.3$ $88.8\pm 0.98$ $79.1\pm 1.2$ $84.2\pm 0.91$ $84.1\pm 0.6$ $88.9\pm 0.42$
ARVGA $68.9\pm 4.07$ $72.6\pm 3.59$ $71.4\pm 1.74$ $75.6\pm 2.23$ $84\pm 0.57$ $87.9\pm 0.48$
SIGVAE $85.1\pm 0.8$ $88.5\pm 0.6$ $77.6\pm 1.58$ $83.4\pm 1.46$
VGNAE $86\pm 0.97$ $90.1\pm 0.55$ $78.7\pm 1.46$ $85\pm 1.23$ $84.6\pm 0.44$ $89.7\pm 0.36$
DVGA $\boldsymbol{97\pm 0.15}$ $\boldsymbol{97.2\pm 0.21}$ $\boldsymbol{94.6\pm 0.48}$ $\boldsymbol{95.3\pm 0.37}$ $\boldsymbol{97.8\pm 0.09}$ $\boldsymbol{98\pm 0.07}$
TABLE III: Results for Link Prediction of citation networks. Higher is better.
Model Cora CiteSeer PubMed
Auc AP Auc AP Auc AP
GAE $91\pm 1.02$ $91.9\pm 0.89$ $90.1\pm 1.28$ $92\pm 0.96$ $96.6\pm 0.06$ $96.9\pm 0.04$
LGAE $93.7\pm 0.38$ $94.5\pm 0.26$ $94.9\pm 0.88$ $95.4\pm 0.9$ $97.6\pm 0.08$ $97.7\pm 0.12$
ARGA $93.1\pm 0.34$ $93.2\pm 0.44$ $93.9\pm 0.63$ $94.3\pm 0.65$ $91.1\pm 1.04$ $90.1\pm 1.08$
GNAE $95\pm 0.34$ $95.6\pm 0.25$ $95.1\pm 0.48$ $96.1\pm 0.4$ $97.5\pm 0.11$ $97.4\pm 0.09$
VGAE $90.9\pm 0.4$ $92.5\pm 0.28$ $89.2\pm 0.99$ $91.3\pm 0.55$ $95.7\pm 0.21$ $96\pm 0.2$
LVGAE $94.7\pm 0.67$ $94.9\pm 0.82$ $94.9\pm 0.36$ $95.4\pm 0.35$ $97.6\pm 0.15$ $97.6\pm 0.23$
ARVGA $91.1\pm 0.7$ $91.2\pm 0.74$ $91.9\pm 1.21$ $92.5\pm 1.13$ $95.4\pm 0.2$ $95.9\pm 0.1$
SIGVAE $91.5\pm 0.68$ $92.6\pm 0.71$ $89.5\pm 0.82$ $91.4\pm 0.57$
VGNAE $95.1\pm 1.27$ $95.6\pm 1.17$ $95.1\pm 0.54$ $96\pm 0.55$ $97.4\pm 0.15$ $97.4\pm 0.14$
@@
TABLE IV: Results for Link Prediction of synthetic graphs with varying numbers of latent factors.(* denotes dataset without features, i.e., $X=\rm{I}$). Higher is better.
@@
6) Experimental Results: Table II and III present the results of link prediction on three real-world citation networks without and with node features, respectively. As observed, DGA and DVGA consistently outperform other GAE-based or VGAE-based methods, which demonstrates the effectiveness of our models’ structure in generating more accurate predictions. For instance, DVGA achieves a prediction accuracy increase of 2.1% and 0.5% when compared to the best-performing VGAE-based baselines on Cora and CiteSeer, respectively. Similarly, DGA demonstrates prediction accuracy improvements of 4.7%, 4.8%, and 1.6% in terms of AUC against the most powerful GAE-Based baselines on Cora, CiteSeer, and PubMed, respectively.
The formation processes of real-world graphs are typically unobservable, which makes it difficult to obtain semantic information about latent factors. In order to gain a deeper understanding and more precise analysis of our models’ behavior, as mentioned earlier before, we generate synthetic datasets. We present link prediction results in Table IV with varying the number of channels ($K$) for latent factors. Our results indicate that our models consistently outperform baselines when varying the number of latent factors. Particularly, as the number of factors increases from 4 to 10, the performance of GAE and VGAE steadily decreases. In contrast, our models’ performance only shows a small decline, indicating that they are not affected by changes in $K$. Especially, DVGA exhibits a substantial improvement over VGAE, surpassing it by 27.3% in AUC and 36.9% in AP on the synthetic graph with ten latent factors. The current methods have limitations in identifying the underlying latent factors that are essential for exploring graph properties, which makes it difficult to learn disentangled representations. In contrast, our approach explicitly takes into account the entanglement of heterogeneous factors and designs a decoder specifically suited for this situation. As a result, our method demonstrates higher accuracy than existing approaches in the task of link prediction across all three datasets.
VI-B Node clustering
In this section, we address the task of unsupervised node clustering in the graph. We utilize the K-means clustering algorithm on the learned node embeddings, setting $K$ according to the number of classes specified in Table I.
TABLE V: Node clustering results on Cora and CiteSeer.
@@ VI-B Node clustering
1) Evaluation metrics: The node labels provided in Table I serve as the true clustering labels for our evaluation. We adopt the approach introduced in [67], where we initially use the Munkres assignment algorithm [68] to match the predicted labels with the true labels. We employ five metrics for validating the clustering results, specifically reporting (i) accuracy (acc); (ii) precision; (iii) F1-score (F1); (iv) normalized mutual information (NMI); (v) average rand index (ARI). Consistent with previous studies such as [23] and [17], our focus is only on Cora and CiteSeer.
2) Experimental Results: Table V shows the results of node clustering for Cora and CiteSeer. The results indicate that both DGA and DVGA exhibit a substantial enhancement across all five metrics in comparison to the remaining baseline methods. For instance, on CiteSeer, DVGA has demonstrated a 49.4% accuracy improvement compared to GNAE and a 45.8% improvement compared to DW. Additionally, it has achieved a 2.5% increase in NMI compared to LVGAE and a 78.8% increase compared to DW. On Cora, DGA has exhibited a 71.6% precision enhancement compared to VGAE and a 28.0% improvement compared to DW. Furthermore, it has shown a 0.6% increase in NMI compared to GNAE and a 27.4% improvement compared to DW. This results further prove the effectiveness and superiority of our models.
TABLE VI: Classification accuracies (in percent). Baseline results from [13].
@@
VI-C Semi-supervised node classification
The node classification task aims to predict the labels of the remaining nodes by utilizing the labels of a subset of nodes in the underlying graph. We adapt DGA and DVGA for node classification on three citation networks by adjusting the loss function to incorporate both the original model loss and a semi-supervised classification term. A hyper-parameter is introduced to control the relative importance of components in the combined objective function. We employ an additional fully connected neural network to map the latent representations to softmax probabilities for the required number of classes. During the optimization process, all parameters are trained jointly.
Experimental Results: The results of node classification are summarized in Table VI. We observe that DVGA outperforms other competing models on Cora and PubMed. DGA also shows comparable performance to CiteSeer, where it closely aligns with the GCN, considered the best baseline. The competitive performance in comparison to state-of-the-art methods highlights the robust capability of our models, even though they are not specifically trained for this task.
VI-D Qualitative Evaluation
For a more comprehensive understanding of our proposed models, we conduct two qualitative experiments for a thorough examination in parallel with VGAE. These evaluations focus on the performance of disentanglement and the informativeness learned in the embeddings.
1) Correlation of Latent Features: We plot the correlation between the dimensions of latent features on the synthetic graph with four latent factors in Figure 3. The figure illustrates the absolute correlation values between elements within the 64-dimensional graph representations obtained from VGAE and DVGA. DVGA’s correlation plot exhibits a clear blockwise correlation pattern, specifically with four distinct diagonal blocks. This suggests that the four channels in DVGA are quite likely to capture mutually exclusive information. Therefore, our method successfully learns disentangled node embeddings and captures the latent factors leading to connections to some degree. This will enhance the interpretability of the link prediction results.
2) Visualization of Node Embeddings:
@@ VI-D Qualitative Evaluation
Figure 4 provides a visual comparison of the learned node embeddings from DW, VGAE, ARVGA, and DVGA on Cora. Utilizing t-SNE [69], we project the node representations into a two-dimensional space. It is evident that DVGA generally learns superior node embeddings, showing high intraclass similarity and interclass differences. These findings emphasize the effectiveness of DVGA in learning disentangled representations of graph-structured data.
VI-E Analysis
TABLE VII: Ablation Studies of DGA.
@@ VI-E Analysis
TABLE VIII: Ablation Studies of DVGA.
@@
1) Ablation Studies: We conduct ablation studies on the essential components of our models to validate their contributions. We compare DVGA with the following three variants: (1) $w/o$ flow: i.e., setting $M=0$. (2) $w/o$ indep: i.e., setting $\lambda=0$. (3) $w/o$ disen: i.e., setting $K=1$, and in this case, our method degenerates into the traditional entangled variational graph auto-encoder. For DGA, as it lacks the flow component in its structure, we consider its similar variants $w/o$ indep and $w/o$ disen. For simplicity, we only report the results of the unsupervised link prediction task, with similar patterns observed in other tasks.
@@
The results for DGA, DVGA, and their variants are presented in Tables VII and VIII. It is evident that removing any module diminishes the predictive capability of the models across all datasets. The elimination of the disentanglement mechanism has had a particularly large impact, resulting in a notable decrease in the performance of the model. For example, the outcomes of DVGA on Cora demonstrate this phenomenon. Therefore, we conclude that each module we designed plays a substantial role in the models’ ability, and their presence collectively enhances the predictive power of the models.
2) Hyper-Parameter Sensitivity: We investigate the sensitivity of three essential hyper-parameters. We present the evaluation results for the channel $K$ on synthetic datasets (Figure 5), as well as the results for all three hyperparameters on Cora (Figure 6), CiteSeer (Figure 7), and PubMed (Figure 8). We present the AUC and AP for link prediction in terms of $K$ for DVGA on a synthetic dataset, along with the AUC for three real-world datasets in terms of all parameters. For the AP, similar results are seen on three real-world datasets, which we have eliminated for the sake of brevity.
@@
  • Number of channels $K$: We initially examine the value of $K$ using the synthetic dataset with eight latent factors. It can be seen that the performance increases with an increment in $K$, reaching its peak at $K=8$, and then decreases. This implies that an optimal $K$ aligning with the ground-truth underlying factors outputs the best results. It also demonstrates that our model learns representations that promote disentanglement and is effective for node-related tasks. Next, we investigate the impact of channel $K$ in three citation networks. As observed previously, the performance exhibits an initial increase, followed by a peak, and then significant declines. This pattern also shows that an appropriate number of channels $K$, if matching the real number of latent factors, can lead to enhanced performance;
@@
  • Regularization coefficient $\lambda$: The regularization coefficient $\lambda$ can affect the model’s performance. A large $\lambda$ tends to excessively emphasize the independence between the latent factors, whereas an excessively small $\lambda$ constrains the influence of the independence regularization. Through empirical observation, we find that setting $\lambda$ to 0.01, 0.1, and 0.001, respectively, leads to satisfactory results for Cora, CiteSeer, and PubMed.
@@
  • Number of disentangle layers $L$: The significance of the number of disentangle layers $L$ lies in the fact that a model with a small $L$ exhibits restricted capacity, potentially hindering its ability to integrate sufficient information from neighbors. From the graphs, we can observe that for Cora, the optimal $L$ is 3; for CiteSeer, it is 5; and for PubMed, it is 7. Additionally, when $L$ is too large, there is a decline in predictive capability, suggesting that excessive message propagation can result in overfitting, thereby compromising model performance;
VI-F Convergence Behavior and Complexity Analysis